# Yutong Xia[Uncaptioned image][Uncaptioned image]  Ao Qu[Uncaptioned image]  Yunhan Zheng[Uncaptioned image]  Yihong Tang[Uncaptioned image]  Dingyi Zhuang[Uncaptioned image] Yuxuan Liang[Uncaptioned image]  Shenhao Wang[Uncaptioned image]  Cathy Wu[Uncaptioned image]  Lijun Sun[Uncaptioned image] Roger Zimmermann[Uncaptioned image]  Jinhua Zhao[Uncaptioned image]

### v2

@@ 4.1 Social Impact
For urban researchers, AutoUrbanCI serves as a powerful assistant in accelerating and refining causal analysis. This allows researchers to focus on higher-level theoretical development and policy interpretations rather than labor-intensive preprocessing and computation. Moreover, AutoUrbanCI facilitates cross-city comparisons and large-scale causal experiments, enabling researchers to analyze urban dynamics across diverse spatial and temporal contexts. This can lead to more robust, generalizable findings that inform evidence-based urban policies and interdisciplinary research collaborations. Additionally, by standardizing causal inference workflows, AutoUrbanCI helps ensure reproducibility and transparency in urban studies, fostering a more rigorous and scalable approach to urban causal research.
@@
AutoUrbanCI provides policymakers with real-time causal analysis, allowing them to evaluate the potential impact of urban policies before implementation. By reducing reliance on correlation-driven decision-making, this automation ensures that interventions, such as traffic regulations, housing policies, or environmental protections, are based on causal insights rather than speculation. This can lead to more effective, data-driven, and proactive governance, ultimately improving urban livability, sustainability, and resource allocation.
@@
In addition, AutoUrbanCI lowers the barriers for citizens, journalists, and grassroots organizations to explore urban issues without requiring deep technical expertise. This fosters greater civic engagement and public awareness of how cities function. Urban studies should not only seek to explain the city but also be accessible to the people who live in it —because cities can only provide something for everybody when they are created by everybody Jacobs (1961). This ensures that data-driven insights contribute to more inclusive and participatory urban decision-making Fox and Wolf (2024).
@@ 4.2 New Human-AI Collaboration Paradigm
Traditionally, AI systems have served as passive tools in urban research workflows, used for predictive modeling Li et al. (2024c), spatial visualization Zhang et al. (2024a), or data preprocessing Liang et al. (2025). However, the introduction of AutoUrbanCI represents a shift toward a more collaborative paradigm. In this framework, AI agents are no longer limited to executing predefined tasks; instead, they actively participate in the scientific reasoning process. Rather than merely accelerating isolated stages of analysis, these agents engage in context-aware interactions that reflect an evolving role: from tool to collaborator. This shift invites deeper questions about how tasks and responsibilities are distributed across human–agent systems, detailed as below.
@@
Moving beyond the traditional pipeline of human-led design, AutoUrbanCI embraces a configurable division of labor between humans and agents. Recent systems Gottweis et al. (2025); Li et al. (2024b); Lu et al. (2024); Swanson et al. (2024) reveal a common design principle: human experts define goals, constraints, and evaluation criteria, while agents handle iterative generation, optimization, and evidence organization. In these systems, natural language interfaces serve as the medium for task specification, correction, and refinement, reinforcing the human’s role as an agenda-setter and critical reviewer. In AutoUrbanCI, this flexible structure allows agents to operate independently for routine tasks such as data preprocessing or method selection, while preserving human judgment for hypothesis formulation, identification strategy, and assumption checking. Multi-agent mechanisms such as self-critique and ranking Lu et al. (2024) can further enhance robustness, but the final authority remains with human users, who validate outputs and determine their scientific and policy relevance. Rather than replacing human reasoning, this structure reinforces it: agents scaffold the causal workflow, while humans remain responsible for framing questions, defining standards, and interpreting outcomes in context.
@@
Taken together, these shifts point toward a broader rethinking of how scientific inquiry is conducted. In AutoUrbanCI, causal analysis is no longer a static pipeline but a collaborative process, structured by agents, guided by human judgment, and mediated through language. As human and machine capabilities become increasingly entangled, urban research must move beyond automation toward shared reasoning systems that are adaptive, interpretable, and participatory.
@@ 4.3 Ethical Considerations
As AI agents take on active roles across the causal workflow, from hypothesis generation to policy-facing outputs, new ethical considerations arise around responsibility and transparency, fairness and bias, and the role of human oversight. First, when automated inferences inform public decisions, it is essential to clarify who is accountable for the assumptions, methods, and consequences. Human-in-the-loop designs can offer oversight, but they must be structured to avoid automation bias and ensure that AI-generated outputs remain contestable. Interpretability should enable not only technical auditing but also meaningful engagement by domain experts and affected communities. Second, urban datasets often encode structural inequities. If agents are trained on dominant narratives or skewed data, they may inadvertently reinforce existing biases. Systems like AutoUrbanCI must embed mechanisms for uncertainty disclosure, sensitivity analysis, and inclusive data practices to mitigate this risk. Third, as AI becomes more embedded in scientific workflows, it raises broader questions around authorship, attribution, and epistemic authority. Rather than replacing human researchers, AutoUrbanCI is designed to support more inclusive, rigorous, and reflexive forms of urban causal reasoning—where domain expertise, civic participation, and algorithmic assistance can co-exist productively.
@@ 5 Summary & the Road Ahead
This Perspective highlights the untapped potential of LLMs to scale urban causal research. Urban causal research plays a vital role in evidence-based policymaking, yet current workflows are often labor-intensive, fragmented, and inaccessible. Our systematic review reveals five key gaps: geographic concentration, overreliance on structured data, limited methods, poor transparency, and slow causal adoption. To address these, we introduces AutoUrbanCI, a modular, LLM-powered framework that supports every stage of the causal pipeline, from hypothesis generation to evaluation, with an example use case. By structuring the process into discrete, automatable tasks, AutoUrbanCI show how AI can help scale causal research while retaining scientific rigor.
@@
AutoUrbanCI is more than a tool. Instead, it is a step toward a more inclusive and participatory paradigm of urban inquiry. It lowers the technical threshold for causal analysis, enabling researchers, policymakers, and civic groups to explore timely urban questions with reduced dependence on coding, manual data handling, or advanced econometrics. In doing so, it fosters greater transparency, reproducibility, and responsiveness in how cities learn and adapt.
@@
Moving forward, we envision three key directions for future work. First, we plan to develop a deployable prototype of AutoUrbanCI to validate its effectiveness in real-world causal research workflows. Second, we aim to refine agent capabilities, especially for handling unstructured and multimodal data and designing rigorous causal inference experiments. Third, we will explore broader deployment scenarios in collaboration with urban researchers and policymakers. We call for collective efforts across AI, urban studies, and broader quantitative social science communities to realize a more inclusive, scalable, and rigorous future for urban causal inference.
@@ Appendix C Methodology for Systematic Review
In Section 2, we present a comprehensive systematic review that maps the current landscape of causal inference in urban studies. Here we detail the methodology used to conduct the review as follows.
Goal.
Our objective was to obtain a quantitative understanding of the distribution of urban causal inference studies across five key dimensions: research topic, geographic focus, data modality, experiment type, and methodological approach.
Step 1: Paper Parsing via LLMs.
We developed a structured prompt to guide GPT through the analysis of each paper, with the aim of extracting relevant information and supporting evidence corresponding to each dimension. To retrieve content from full-text PDFs, we utilized the OpenAI API’s file upload feature [^5] https://platform.openai.com/docs/guides/pdf-files?api-mode=chat, which allows direct querying over document contents. This step produced initial annotations per dimension for every paper.
Step 2: Iterative Label Refinement.
For each dimension, we designed an iterative clustering pipeline that leverages GPT-4o to semantically group similar entries and propose representative labels. The algorithm identifies recurring patterns, merges equivalent concepts, and refines the terminology. This step outputs a set of standardized keywords describing each paper across the five dimensions.
Step 3: Human Validation.
To ensure accuracy and domain relevance, we manually reviewed and refined the final set of clusters and labels using expert judgment. This final validated dataset served as the basis for the visualizations and quantitative analyses presented in the main text.

### v3

@@ 4.1 Social Impact
For urban researchers, UrbanCIA serves as a powerful assistant in accelerating and refining causal analysis. This allows researchers to focus on higher-level theoretical development and policy interpretations rather than labor-intensive preprocessing and computation. Moreover, UrbanCIA facilitates cross-city comparisons and large-scale causal experiments, enabling researchers to analyze urban dynamics across diverse spatial and temporal contexts. This can lead to more robust, generalizable findings that inform evidence-based urban policies and interdisciplinary research collaborations. Additionally, by standardizing causal inference workflows, UrbanCIA helps ensure reproducibility and transparency in urban studies, fostering a more rigorous and scalable approach to urban causal research.
@@
UrbanCIA provides policymakers with real-time causal analysis, allowing them to evaluate the potential impact of urban policies before implementation. By reducing reliance on correlation-driven decision-making, this automation ensures that interventions, such as traffic regulations, housing policies, or environmental protections, are based on causal insights rather than speculation. This can lead to more effective, data-driven, and proactive governance, ultimately improving urban livability, sustainability, and resource allocation.
@@
In addition, UrbanCIA lowers the barriers for citizens, journalists, and grassroots organizations to explore urban issues without requiring deep technical expertise. This fosters greater civic engagement and public awareness of how cities function. Urban studies should not only seek to explain the city but also be accessible to the people who live in it —because cities can only provide something for everybody when they are created by everybody Jacobs (1961). This ensures that data-driven insights contribute to more inclusive and participatory urban decision-making Fox and Wolf (2024).
@@ 4.2 New Human-AI Collaboration Paradigm
Traditionally, AI systems have served as passive tools in urban research workflows, used for predictive modeling Li et al. (2024c), spatial visualization Zhang et al. (2024a), or data preprocessing Liang et al. (2025). However, the introduction of UrbanCIA represents a shift toward a more collaborative paradigm. In this framework, AI agents are no longer limited to executing predefined tasks; instead, they actively participate in the scientific reasoning process. Rather than merely accelerating isolated stages of analysis, these agents engage in context-aware interactions that reflect an evolving role: from tool to collaborator. This shift invites deeper questions about how tasks and responsibilities are distributed across human–agent systems, detailed as below.
@@
Moving beyond the traditional pipeline of human-led design, UrbanCIA embraces a configurable division of labor between humans and agents. Recent systems Gottweis et al. (2025); Li et al. (2024b); Lu et al. (2024); Swanson et al. (2024) reveal a common design principle: human experts define goals, constraints, and evaluation criteria, while agents handle iterative generation, optimization, and evidence organization. In these systems, natural language interfaces serve as the medium for task specification, correction, and refinement, reinforcing the human’s role as an agenda-setter and critical reviewer. In UrbanCIA, this flexible structure allows agents to operate independently for routine tasks such as data preprocessing or method selection, while preserving human judgment for hypothesis formulation, identification strategy, and assumption checking. Multi-agent mechanisms such as self-critique and ranking Lu et al. (2024) can further enhance robustness, but the final authority remains with human users, who validate outputs and determine their scientific and policy relevance. Rather than replacing human reasoning, this structure reinforces it: agents scaffold the causal workflow, while humans remain responsible for framing questions, defining standards, and interpreting outcomes in context.
@@
Taken together, these shifts point toward a broader rethinking of how scientific inquiry is conducted. In UrbanCIA, causal analysis is no longer a static pipeline but a collaborative process, structured by agents, guided by human judgment, and mediated through language. As human and machine capabilities become increasingly entangled, urban research must move beyond automation toward shared reasoning systems that are adaptive, interpretable, and participatory.
@@ 4.3 Ethical Considerations
As AI agents take on active roles across the causal workflow, from hypothesis generation to policy-facing outputs, new ethical considerations arise around responsibility and transparency, fairness and bias, and the role of human oversight. First, when automated inferences inform public decisions, it is essential to clarify who is accountable for the assumptions, methods, and consequences. Human-in-the-loop designs can offer oversight, but they must be structured to avoid automation bias and ensure that AI-generated outputs remain contestable. To prevent misuse of AI tools by powerful corporations or institutions for their own advantage, interpretability should enable not only technical auditing but also meaningful engagement by domain experts and affected communities. Second, urban datasets often encode structural inequities. If agents are trained on dominant narratives or skewed data, they may inadvertently reinforce existing biases. Systems like UrbanCIA must embed mechanisms for uncertainty disclosure, sensitivity analysis, and inclusive data practices to mitigate this risk. Third, as AI becomes more embedded in scientific workflows, it raises broader questions around authorship, attribution, and epistemic authority. Rather than replacing human researchers, UrbanCIA is designed to support more inclusive, rigorous, and reflexive forms of urban causal reasoning—where domain expertise, civic participation, and algorithmic assistance can co-exist productively.
@@ 5 Summary & the Road Ahead
This Perspective highlights the untapped potential of LLMs to scale urban causal research. Urban causal research plays a vital role in evidence-based policymaking, yet current workflows are often labor-intensive, fragmented, and inaccessible. Our systematic review reveals five key gaps: geographic concentration, overreliance on structured data, limited methods, poor transparency, and slow causal adoption. To address these, we introduces UrbanCIA, a modular, LLM-powered framework that supports every stage of the causal pipeline, from hypothesis generation to evaluation, with an example use case. By structuring the process into discrete, automatable tasks, UrbanCIA show how AI can help scale causal research while retaining scientific rigor.
@@
UrbanCIA is more than a tool. Instead, it is a step toward a more inclusive and participatory paradigm of urban inquiry. It lowers the technical threshold for causal analysis, enabling researchers, policymakers, and civic groups to explore timely urban questions with reduced dependence on coding, manual data handling, or advanced econometrics. In doing so, it fosters greater transparency, reproducibility, and responsiveness in how cities learn and adapt.
@@
Moving forward, we envision three key directions for future work. First, we plan to develop a deployable prototype of UrbanCIA to validate its effectiveness in real-world causal research workflows. Second, we aim to refine agent capabilities, especially for handling unstructured and multimodal data and designing rigorous causal inference experiments. Third, we will explore broader deployment scenarios in collaboration with urban researchers and policymakers. We call for collective efforts across AI, urban studies, and broader quantitative social science communities to realize a more inclusive, scalable, and rigorous future for urban causal inference.
@@ Appendix C Methodology for Systematic Review
In Section 2, we present a comprehensive systematic review that maps the current landscape of causal inference in urban studies. The objective is to obtain a quantitative understanding of the distribution of urban causal inference studies across five key dimensions: research topic, geographic focus, data modality, experiment type, and methodological approach. Here we detail the methodology used to conduct the review as follows. The prompts used for each step and the code implementing this procedure can be found at: https://github.com/quao627/Urban-CI-LitReview.
Step 1: Identify Papers That Conduct Causal Inference Studies
Among the 2,428 papers downloaded from the Cities journal, we identify those conducting causal inference studies using a crafted prompt with the GPT-4o model. This filtering process involves two steps. First, we exclude papers that are clearly out of scope by examining only their titles and abstracts. Second, the LLM accesses the full text of the remaining papers and labels each one accordingly. This two-step process helps reduce costs by eliminating irrelevant papers early using a low-cost method.
Step 2: Paper Parsing via LLMs.
We develop a structured prompt to guide GPT through the analysis of each paper, with the aim of extracting relevant information and supporting evidence corresponding to each dimension. To retrieve content from full-text PDFs, we utilize the OpenAI API’s file upload feature [^5] https://platform.openai.com/docs/guides/pdf-files?api-mode=chat, which allows direct querying over document contents. This step produced initial annotations per dimension for every paper.
Step 3: Iterative Label Refinement.
For each dimension, we design an iterative clustering pipeline that leverages GPT-4o to semantically group similar entries and propose representative labels. The algorithm identifies recurring patterns, merges equivalent concepts, and refines the terminology. This step outputs a set of standardized keywords describing each paper across the five dimensions.
Step 4: Human Validation.
To ensure accuracy and domain relevance, we manually review and refine the final set of clusters and labels using expert judgment. This final validated dataset serves as the basis for the visualizations and quantitative analyses presented in the main text.